OpenStudy (anonymous):

7/10 of the body is water Explain how you can use this information to determine how many pounds of water are in a person weighing 180 pounds How many pounds of water are in a person weighing 180 pounds How many pounds of a 200 pound person are NOT water if you have any idea on how to help... telll me and explain! thxx

OpenStudy (anonymous):

Ok so given that 7/10 of weight = water, use the equation: First answer: 7/10 * Weight = Water weight Second answer: 180 * 7/10 = 126 lb Third answer: 3/10 * Weight = Non-water Weight 3/10 * 200 = 60 lb

OpenStudy (anonymous):

You are given: 7/10 of the body is water 7/10 of the "body" translates to 7/10 of the "weight" since you are trying to find the various weights of water and non-water for differently weighing bodies. So 7/10 of the weight is the water weight. So that's where I get the numerical equation of: 7/10 * Weight = Water weight For the second answer, you want to use that equation when the weight is 180 pounds. So plug in 180 into Weight. This gives you 7/10 * 180 = Water weight = 126 lb Third answer: I used the assumption that any weight that is non-water is therefore the weight that is not water. So since 7/10 of the body weight is water, 3/10 of the body weight is non-water. This gives the equation of: 3/10 * Weight = Non-water weight Plug in the weight of 200 lbs and you get 3/10 * 200 = 60 lb.